OpenStudy (anonymous):

Lim (x+1)^2-1/x ? x>>>0

OpenStudy (agent47):

did u mean: (x+1)^2 - (1/x) ?

OpenStudy (anonymous):

no i mean divided by x

OpenStudy (agent47):

oh the whole thing?

OpenStudy (anonymous):

yes

OpenStudy (agent47):

((x+1)^2-1)/x

OpenStudy (agent47):

Ok, (x^2+2x+1-1)/x = (x^2+2x)/x=x(x+2)/x=x+2

OpenStudy (agent47):

Now you can simply plug in x=0, and get the limit: 0+2=2

OpenStudy (anonymous):

lim(x+1)^2 -1/x

OpenStudy (anonymous):

where x>>>1

OpenStudy (anonymous):